OpenStudy (anonymous):

What is the 6th term of the geometric sequence where a1 = 625 and a3 = 25? 0.0016 0.008 0.04 0.2

OpenStudy (anonymous):

It is 0.2 you divide by 5 to get the next term 625/5=125 125/5=25 25/5=5 5/5=1 1/5=.2

OpenStudy (nurali):

nth term = a r ^ ( n - 1 ) 3rd term = 25 = 625 r ^ ( 3 - 1 ) 25 = 625 r ² r^2 = 25 / 625 r = 5 / 25 r= 1 / 5 6th term = 625 ( 1 / 5 ) ^ ( 6 - 1 ) = 625 ( 1 / 5 ) ^5 = 625 ( 1 / 3,125 ) = 1 / 5 = 0.2

OpenStudy (anonymous):

Thank you!

OpenStudy (nurali):

My Pleasure.